OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue).

OpenStudy (kropot72):

\[P(red)=\frac{4}{9}\] \[P(blue)=\frac{3}{9}\] \[P(red\ and\ blue)=\frac{4}{9}\times \frac{3}{9}\]